OpenStudy (aripotta):

In an election for senior class president, Alice received 30 votes more than David. If the ratio of the number of votes Alice received to the number of votes David received was 8 to 7, how many votes did David receive? a. 30 b. 49 c. 64 d. 210 e. 240

OpenStudy (anonymous):

You'll have 1 equation, so let x = votes Aliced received and y = voted David recieved \[\frac{ x }{ y } = \frac{ 8 }{ 7 }\]

OpenStudy (kropot72):

Let David's votes be d votes. Then Alice's votes = d + 30. So we can write this equation: \[\large \frac{d+30}{d}=\frac{8}{7}\] Now just solve for d.

OpenStudy (anonymous):

That works nicely to I believe, and 2 equations* letting x = y + 30

OpenStudy (kropot72):

Quite right :)

OpenStudy (aripotta):

7(d + 30) = 8d 7d + 210 = 8d 210 = d oh lol. easier than i thought :) thank you. i thought i would have to make the proportion 8/15, not 8/7

OpenStudy (kropot72):

You're welcome :)